What is the answer to -7-(-8)
Found 2 solutions by Alan3354, tommyt3rd:
Answer by Alan3354(69443)

What is the answer to -7-(-8)
= -7 + 8

Answer by tommyt3rd(5050)

-7-(-8)=
-7+8=1